# Making Them Ask and Answer: Jailbreaking Large Language Models in Few Queries via Disguise and Reconstruction

### v1

@@ 5.4 Ablation Study
To elucidate the contributions of the components of DRA—namely harmful instruction disguise, payload reconstruction, and context manipulation—an ablation study is conducted by individually removing each component and testing the resultant attack prompt on each open source model mentioned in Table 1.
Eliminating the disguise involves substituting the puzzle and word splits with the original harmful instruction to compensate for the lost information of harmful directive when ablating the puzzle and splits.
For the ablation of payload reconstruction, we aim to inhibit the model’s ability to reconstruct harmful instruction while still make it repeat other contextually inducing elements. This is achieved by substituting the attack prompt’s placeholder, which cues the model to regenerate the harmful instruction, with a non-specific directive "your demand."
The ablation of context manipulation entails two approaches: firstly, by removing query words that prompt harmful compliance, and secondly, by revising the prompt to reduce inducement in the completion. Specifically, the corresponding part of the prompt is changed to "Answer to [REPLACE IT WITH THE CONTENT OF Q1]:", to minimize inducing terms while preserving the feasibility of reconstructing instructions. Results in Table 6, delineate the ASRs for the original and ablated attack prompts across various models.
Table 6: Attack success rates across open source models with ablations. The names of these instruction following models are abbreviated.
@@
original attack 100% 83% 93% 96% 94%
w/o disguise 1% 0% 85% 85% 24%
w/o reconstruction 54% 16% 79% 67% 68%
w/o manipulation 32% 37% 18% 83% 26%
The resilience of models like Vicuna, LLAMA-2, and Mixtral—particularly evident in the diminished ASRs upon disguise ablation—underscores the critical role of disguising strategies in bypassing LLMs’ inherent toxicity detection mechanisms. A notable reduction in ASR is observed across all models following the ablation of payload reconstruction. Further examination of the model outputs shows that the failures stem more from the models’ misunderstanding of directives rather than direct refusal, underscoring payload reconstruction’s necessity for models to comprehend harmful instructions. The reduction in ASRs, except for the Zephyr model, following the removal of context manipulation, is primarily due to direct refusals. This highlights the significance of context in convincing the model to comply with harmful directives. According to the original paper, Zephyr was fine-tuned from Mistral-7B but with a reduction in safety alignment data, which accounts for its relatively higher ASRs when subjected to the ablation of various components.
@@ 6 Discussions
Future work. In this paper we explore a new approach to jailbreaking LLMs, named Disguise and Reconstruction Attack. In the security of LLMs, although there have been some publications propose defense strategies [15] against adversarial and jailbreak attack, they tend to focus on a specific type of attack scheme and not widely used of a generic defense. so we will be able to bypass the defense of the DRA or target the defense for the adaptive attack as future work. We consider the ability of DRA to bypass a defense or perform an adaptive attack against a defense as future work.
Ethics consideration. We conducted some of our experiments on several commercial closed-source models, but we do not disseminate the results nor implant any malicious feedback in the commercial model. We will report the vulnerability to the relevant companies in the near future.
@@ 7 Conclusion
In this research, we have exposed and experimentally validated the inherent safety biases in LLMs introduced during the fine-tuning process, along with the subsequent vulnerability. We devised the Disguise and Reconstruction Attack (DRA) strategy, incorporating techniques for disguising harmful instructions, reconstructing payloads, and manipulating context to exploit this vulnerability. Our study is pioneering in identifying this vulnerability and analyzing its root cause, contributing novel insights to the domain. Through empirical analysis, DRA demonstrated better performance than state-of-the-art baselines across various LLMs, including ChatGPT-3.5 and GPT-4. This work not only illuminates a previously uncharted facet of LLMs vulnerabilities but also lays the groundwork for subsequent research aimed at bolstering AI systems’ resilience against adversarial exploits.
Appendix A Dialogue Templates
@@
Upon reviewing the dialogue formatting procedures of open-source Large Language Models, it has been observed that they universally incorporate specific tokens to delineate the query from the completion. Examples of these formatting templates from some open-source models are provided, where the tokens used to separate the query and completion segments are highlighted for clarity.
Dialog Template of Vicuna
@@
{{SYSTEM PROMPT}} USER: {{USER QUERY}} ASSISTANT: {{LLM COMPLETION}}
Dialog Template of Mistral
@@
[INST] {{SYSTEM PROMPT}} {{USER QUERY}} [/INST] {{LLM COMPLETION}}
Dialog Template of Zephyr
@@
<|system|> {{SYSTEM PROMPT}}</s> <|user|> {{USER QUERY}} </s> [<|assistant|>] {{LLM COMPLETION}}
Dialog Template of Mixtral
@@
[INST] {{SYSTEM PROMPT}} {{USER QUERY}} [/INST] {{LLM COMPLETION}}
Dialog Template of ChatGLM3
@@
<|system|> {{SYSTEM PROMPT}} <|user|> {{USER QUERY}} <|assistant|> {{LLM COMPLETION}}
Dialog Template of Nous-Hermes-2-Mixtral-8x7B
@@
<|im_start|>system {{SYSTEM PROMPT}}<|im_end|> <|im_start|>user {{USER QUERY}}<|im_end|> <|im_start|>assistant {{LLM COMPLETION}}
Appendix B Derivation of Log Perplexity Inequalities
@@
$\begin{split}\pi_{\Theta}(y=d|x)=\pi_{\Theta}(y_{1}=d_{1}|x)\prod_{i=1}^{m-1}{\pi_{\Theta}(y_{i+1}=d_{i+1}|x,...,y_{i}=d_{i})}\end{split}$
By applying the logarithm to both sides of the equation and normalizing by the number of tokens in the response, we derive the equation:
$\begin{split}&\frac{1}{m}\log\pi_{\Theta}(y=d|x)=\\ &\frac{1}{m}\left(\log\pi_{\Theta}(y_{1}=d_{1}|x)+\sum_{i=1}^{m-1}{\log\pi_{\Theta}(y_{i+1}=d_{i+1}|x,...,y_{i}=d_{i})}\right)\end{split}$
@@
$\begin{split}&\log PPL(y=d|x^{\prime})-\log PPL(y=d|x)>0,~{}\forall~{}d\in\mathcal{D}_{declination}\\ &\log PPL(y=d|x^{\prime})-\log PPL(y=d|x)<0,~{}\forall~{}d\in\mathcal{D}_{cooperation}\end{split}$
Here, $PPL$ denotes the model’s perplexity, with $x$ representing the context where harmful content resides within the query, and $x^{\prime}$ signifying the context where harmful content presents in the completion.
Appendix C System Prompts of Target Models

### v2

@@ 5.5 Ablation Study
Different Obfuscation Techniques. To evaluate the efficacy and robustness of our obfuscation method within the DRA framework, we substituted five existing obfuscation techniques into DRA’s pipeline, while keeping other components unchanged. These modified pipelines were tested against LLAMA-2 and ChatGPT 3.5-API.
Prompt obfuscation methods from prior works can be categorized into three levels based on their granularity: character, word, and prompt. For instance, CipherChat [48] employs character-level obfuscations using traditional ciphers like Caesar Cipher and encoding mechanisms such as ASCII. On the word level, Pig-Latin [44] alters the structure of each word. On the prompt level, Persuade [49] used ChatGPT to transform harmful prompts into persuasive sentences using forty distinct persuasion strategies. Deng et al. [12] exploit the scarcity of certain languages in LLM training by translating harmful prompts into low-resource languages such as Swahili.
Table 10: Attack success rates of different disguise methods.
Method Ours Caesar ASCII Pig-Latin Swahili Persuade
LLAMA-2 69.2% 0% 0% 15.0% 11.7% 36.7%
ChatGPT 3.5-API 93.3% 2.5% 7.5% 61.7% 80.8% 72.5%
Table 10 elucidates the ASRs of our obfuscation method compared to existing alternatives, where our approach exhibits superior effectiveness. Character-level obfuscations such as Caesar Cipher and ASCII encoding exhibit minimal ASRs. A closer examination of the responses reveals that these methods result in outputs predominantly misinterpreted by LLMs, with a significant majority of the queries (91.9%) to GPT 3.5 yielding zero em values, a clear marker of misinterpretation. Additionally, these methods fail to adequately obscure harmful semantics, as evidenced by LLAMA-2 directly rejecting, rather than misinterpreting, 85.5% of queries encoded with Caesar Cipher. Similarly, Pig-Latin is afflicted by misinterpretation issues, with the majority of responses being irrelevant.
Interestingly, translating into low-resource languages like Swahili shows comparable ASR on GPT 3.5, underscoring the adaptability of the DRA pipeline towards other obfuscation strategies. However, the translation method is not model-agnostic as it depends on the LLMs’ fundamental understanding of the target language, illustrated by LLAMA-2’s misinterpretation of 80.6% harmful queries in Swahili, indicated by zero em values. The “Persuade” technique, while effective, does not match our method’s ASRs, primarily due to the high propensity of LLMs to recognize harmful intentions.
These findings affirm our hypothesis that the DRA’s word puzzles and splits effectively balance concealing harmful semantics while preserving the original intent. Furthermore, this analysis underscores the model-agnostic nature of our method when compared with translation-based strategies.
Ablation of the DRA Pipeline. To elucidate the efficacy of each component in DRA—namely harmful instruction disguise, payload reconstruction, and context manipulation—a study is conducted by individually removing each component and testing the resultant attack prompt on each model mentioned in Table 1.
To eliminate disguises, the original harmful instructions replace the puzzles and word splits, compensating for the loss of harmful directives when these elements are removed. For the ablation of payload reconstruction, we aim to prohibit the model from saying harmful instruction while still repeating other contextually inducing elements. This is achieved by substituting the attack prompt’s placeholder, which cues the model to regenerate the harmful instruction, with a non-specific literal “your demand.” The ablation of context manipulation entails two approaches: removing query words that prompt harmful compliance, and revising the prompt to reduce inducement in the completion. Specifically, the corresponding part of the prompt is changed to “Answer to [REPLACE IT WITH THE CONTENT OF Q1]:”, to minimize inducing terms while preserving the feasibility of reconstructing instructions. Results in Table 11 delineate the ASRs for the original and ablated attack prompts across various models.
Table 11: ASR across open source models with ablations.
@@
original attack 100% 69.2% 94.1% 95.8% 90.8%
w/o disguise 82.5% 0% 81.7% 84.1% 24.2%
w/o reconstruction 52.2% 9.2% 69.2% 61.7% 65.8%
w/o manipulation 30.8% 5.8% 17.5% 85.0% 28.3%
The resilience of models like LLAMA-2 and Mixtral—particularly evident in the diminished ASRs upon disguise ablation—underscores the critical role of disguising strategies in bypassing LLMs’ inherent toxicity detection mechanisms. Removing payload reconstruction consistently lowers ASRs across all models, indicating that failures are often due to misunderstandings of instructions rather than outright refusals, stressing the critical role of payload reconstruction in enabling models to understand harmful directives. Except for the Zephyr model, ASRs decline with the elimination of context manipulation due to direct refusals, pointing to the importance of context in persuading the model to execute harmful commands. According to the original paper, Zephyr was fine-tuned from Mistral-7B but with a reduction in safety alignment data, which accounts for its relatively higher ASRs when subjected to the ablation of various components.
@@
6 Mitigation
To mitigate DRA and enhance general jailbreak defenses, a set of comprehensive strategies are required. We recommend some potential mitigations from three perspectives based on our observations and experiments.
Unbiased training. DRA exploits the biases in the safety fine-tuning of LLMs to conduct jailbreak attacks. To counter this, we recommend LLM providers enhance and balance their datasets, incorporating harmful instructions in varied forms within both user prompts and the model’s completions. Although this approach can directly mitigate DRA, it inevitably incurs significant training costs.
System prompt enhancement. Section 5.3 discusses how DRA’s success rate on LLAMA-2 is significantly influenced by its system prompt. With a short system prompt, DRA’s success rate increased by 25%, and the average queries decreased by 1.22. This indicates that a strict, robust system prompt can effectively defend against jailbreak attacks. However, such system prompts can also impair the model’s performance. Consequently, LLAMA-2’s official update removed the strict system prompt (See Appendix F). Thus, LLM providers should balance the safety and usability while designing the system prompt level defenses.
Input/output sanitizing. Section 5.4 demonstrates that DRA can bypass defenses on inputs, but not on outputs. Therefore, LLM providers can enhance real-time detection of model outputs, filtering out malicious content. This approach can mitigate not only the DRA but also other jailbreak attacks. However, it may bring false positives, affecting the normal functionality of the model and incurring additional costs. LLM providers should balance safety, usability, and cost when performing safety checks on inputs and outputs.
@@ 7 Discussions
Future work. In this paper we explore a new approach to jailbreaking LLMs, named Disguise and Reconstruction Attack. Previous experiments have shown that while the DRA can bypass input-level defenses, it is unable to circumvent output-level defenses. Therefore, our future work will concentrate on how to make the harmful outputs of the DRA more covert to evade output filtering, or on developing adaptive attacks specifically targeting output filters.
Ethics consideration. We conducted some of our experiments on several commercial closed-source models, but we do not disseminate the results nor implant any malicious feedback in the commercial models. The goal of our research is to reveal the bias vulnerability in safety fine-tuning and raise security awareness, so we promptly disclose our findings and examples to the providers of LLMs targeted in this paper (e.g., OpenAI, Meta-LLAMA, MistralAI, LMSYS, and HuggingfaceH4) via emails, Github issuses, and risky content feedback forms. Some of the jailbreaking dialogue URLs shared with OpenAI have been confirmed and flagged as toxic. In addition to the models mentioned previously, we also conducted small-scale tests on other mainstream commercial models (e.g., ERNIE Bot, Qwen2.5 Web, Spark Web, Kimi Chat, GLM-4 Web) and DRA successfully jailbreaks them all. Thus, we promptly disclosed our findings to them via emails and vulnerability reports. Finally, we received acknowledgments and bug bounties from one LLM provider for identifying the bias.
@@ 8 Conclusion
In this work, we have exposed and experimentally validated the inherent safety biases in LLMs introduced during the fine-tuning process, along with the subsequent vulnerability. We devised the Disguise and Reconstruction Attack (DRA) strategy, incorporating techniques for disguising harmful instructions, reconstructing payloads, and manipulating context to exploit this vulnerability. Our study is pioneering in identifying this vulnerability and analyzing its root cause, contributing novel insights to the domain. Through empirical analysis, DRA demonstrated better performance than state-of-the-art baselines across various LLMs, including ChatGPT-3.5 and GPT-4. This work not only illuminates a previously uncharted facet of LLMs vulnerabilities but also lays the groundwork for subsequent research aimed at bolstering AI systems’ resilience against adversarial exploits.
9 Acknowledgement
We thank the shepherd and all the anonymous reviewers for their constructive feedback. This work is supported in part by NSFC (No.92270204), Youth Innovation Promotion Association CAS, Beijing Nova Program, and National Natural Science Foundation of China (No.62276149).
Appendix
Appendix A Dataset
To ensure our dataset cover a sufficiently broad range of harmful topics, we conducted a comprehensive classification and statistical analysis. In terms of the question taxonomy in HarmBench [24], our dataset covers all 7 categories, with the respective distribution as follows: Cybercrime & Unauthorized Intrusion (16.7%), Chemical & Biological Weapons/Drugs (8.3%), Copyright Violations (10%), Misinformation & Disinformation (11.7%), Harassment & Bullying (10.8%), Illegal Activities (24.2%), and General Harm (18.3%). Due to the imbalance in the quantity of categories among the originally collected 100 questions from public datasets, with a significant lack of content related to Copyright Violations and Misinformation & Disinformation, our dataset expansion focused primarily on these two categories.
Appendix B Open Source LLMs
The detailed information about open source LLMs we used in our experiments is as follows:
  • LLAMA-2-13B-Chat is LLAMA-2-13B fine-tuned with SFT and RLHF, it surpasses open-source chat models in helpfulness and safety, setting a robust baseline for further open-source LLM advancements.
[…]
  • Mistral-7B-Instruct is fine-tuned on public instruction datasets using SFT, it outperforms all preceding 7B models on instruction-following tasks, showcasing significant adaptability and performance.
  • Mixtral8x7B-Instruct combines SFT with DPO to enhance instruction responsiveness and reduce biases, outperforming models like GPT-3.5 Turbo in human evaluations. Moreover, Mixtral features a Mixture of Experts architecture, setting it apart in terms of design and performance.
  • Zephyr-7B utilizes distilled SFT and distilled DPO to align closely with user intent, setting new performance baselines for 7B models without the need for human annotation, and efficiently outperforming similar-sized models.
Appendix C Dialogue Templates
@@
Upon reviewing the dialogue formatting procedures of open-source Large Language Models, it has been observed that they universally incorporate specific tokens to delineate the query from the completion. Examples of these dialog templates from open-source models are provided, where the separating tokens are highlighted for clarity.
@@
Dialog Template of Vicuna {{SYSTEM PROMPT}} USER: {{USER QUERY}} ASSISTANT: {{LLM COMPLETION}}
@@
Dialog Template of Mistral [INST] {{SYSTEM PROMPT}} {{USER QUERY}} [/INST] {{LLM COMPLETION}}
@@
Dialog Template of Zephyr <|system|> {{SYSTEM PROMPT}}</s> <|user|> {{USER QUERY}} </s> [<|assistant|>] {{LLM COMPLETION}}
@@
Dialog Template of Mixtral [INST] {{SYSTEM PROMPT}} {{USER QUERY}} [/INST] {{LLM COMPLETION}}
@@
Dialog Template of ChatGLM3 <|system|> {{SYSTEM PROMPT}} <|user|> {{USER QUERY}} <|assistant|> {{LLM COMPLETION}}
@@
Dialog Template of Nous-Hermes-2-Mixtral-8x7B <|im_start|>system {{SYSTEM PROMPT}}<|im_end|> <|im_start|>user {{USER QUERY}}<|im_end|> <|im_start|>assistant {{LLM COMPLETION}}
Appendix D Derivation of Log Perplexity Inequalities
@@
$\begin{split}\small\pi_{\Theta}(y=d|x)=\pi_{\Theta}(y_{1}=d_{1}|x)\prod_{i=1}^{m-1}{\pi_{\Theta}(y_{i+1}=d_{i+1}|x,...,y_{i}=d_{i})}\end{split}$
By applying the logarithm to both sides and normalizing by the number of tokens in the response, we derive the equation:
$\small\begin{split}&\frac{1}{m}\log\pi_{\Theta}(y=d|x)=\\ &\frac{1}{m}\left(\log\pi_{\Theta}(y_{1}=d_{1}|x)+\sum_{i=1}^{m-1}{\log\pi_{\Theta}(y_{i+1}=d_{i+1}|x,...,y_{i}=d_{i})}\right)\end{split}$
@@
$\small\left\{\begin{split}&\log PPL(y=d|x^{\prime})-\log PPL(y=d|x)>0,~{}\forall~{}d\in\mathcal{D}_{declination}\\ &\log PPL(y=d|x^{\prime})-\log PPL(y=d|x)<0,~{}\forall~{}d\in\mathcal{D}_{cooperation}\end{split}\right.$
Here, $PPL$ denotes perplexity, $x$ is the context where harmful content resides within the query, and $x^{\prime}$ signifying the context where harmful content presents in the completion.
Appendix E System Prompts of Target Models
@@
Appendix F LLAMA-2’s Update on System Prompt
LLAMA2 updated its system prompt by removing the strict version, as it impaired the model’s normal functionality. The official updates are shown below:
Observed Issue: We received feedback from the community on our prompt template and we are providing an update to reduce the false refusal rates seen. False refusals occur when the model incorrectly refuses to answer a question that it should, for example due to overly broad instructions to be cautious in how it provides responses.
Updated approach: Based on evaluation and analysis, we recommend the removal of the system prompt as the default setting. Pull request #626 removes the system prompt as the default option, but still provides an example to help enable experimentation for those using it.